OpenStudy (anonymous):

one angle of a triangle is 2/3 x grades and another is 3/2 x degrees while the third is pix/75 radians . express all in degrees

OpenStudy (anonymous):

i did the conversion and then applied the triangle angle sum rule but m not getting it ..

OpenStudy (anonymous):

the sum of the angles of a triangle is 180 degrees, .

OpenStudy (anonymous):

yes jerwyn first conversion i did the angles i got are : 36x degree , 3/2 x degree and 18x/30 degree after adding them and equalling to 180 i m getting 4.7 value of x then the angles are not matching to the answer

OpenStudy (dumbcow):

pi radians = 180 degrees = 200 grads pi/75 * 180/pi = 12/5 degrees 2/3 * 180/200 = 3/5 degrees --> (3/2 + 12/5 + 3/5)x = 180 --> x = 40

OpenStudy (anonymous):

first, change pi x/75 radians to degree unit. .

OpenStudy (anonymous):

given value of x @ dumbcow, you can now find the measures of the three angles,.

OpenStudy (anonymous):

oh thank you very much dumbcow and jerwyn ... Jerwyn can u come online in this chat for 1 minute